OpenStudy (anonymous):

Evaluate the limit below involving absolute value: lim x→1+ (x^2 + 4x − 5) / (|x − x^2|)

OpenStudy (dumbcow):

Factor \[\rightarrow \frac{(x-1)(x+5)}{|x(1-x)|}\] then split absolute value into 2 cases...since you are evaluating at 1 from right x is positive and (1-x) will be neg \[\rightarrow \frac{(x-1)(x+5)}{-x(1-x)} = \frac{(x-1)(x+5)}{x(x-1)} = \frac{x+5}{x}\] evaluate at 1 \[\lim = 6\]

OpenStudy (anonymous):

since i'm evaluating 1 from the right, why is denominator -x(1-x) instead of (x(1-x))

OpenStudy (dumbcow):

because (1-x) is neg when x>1 ....the abs value makes it pos, or in other words multiply by -1

OpenStudy (dumbcow):

its a way to get rid of the abs value signs without affecting the value

OpenStudy (anonymous):

whenever i have an absolute, do i always factor and focus on the inside parentheses

OpenStudy (anonymous):

in situations like this?

OpenStudy (dumbcow):

yes...you have to pay attention to what values of x you are using and if it makes the inside "neg" or not

OpenStudy (dumbcow):

if its all pos values then drop the abs value signs and don't change anything

OpenStudy (anonymous):

oh ok. now i see. thank you so much!

OpenStudy (dumbcow):

yw ...for practice try to see what the limit is if you evaluate from the left

OpenStudy (anonymous):

hmm.. this doesn't seem right but i came up with 6. from the left, it should be negative. so when x<1 the x(x-1) should do

OpenStudy (dumbcow):

you are right, it should be -6 lets see...when x<1 , (1-x) is positive so dont change anything \[\frac{(x-1)(x+5)}{x(1-x)} = \frac{(x-1)(x+5)}{-x(x-1)} = \frac{x+5}{-x} = -6\] this time you factor out a neg, so you can cancel the (x-1)

OpenStudy (anonymous):

why is it (x-1)? wouldn;'t it be negative. wouldn't the denominator be -x(1-x)

OpenStudy (dumbcow):

no but then you have exactly what we had before the original abs value was |x(1-x)| , the inside can be either + or - depending on x right so if x(1-x) is "-" then put a neg sign in front to make it pos if x(1-x) is "+" then leave it alone

OpenStudy (anonymous):

could it be x(1-x)

OpenStudy (dumbcow):

you mean could \[|x(1-x)| = x(1-x)\] ? yes if 0<x<1

OpenStudy (anonymous):

oh i thought it could work for 1- since the inside of absolute value is negative

OpenStudy (dumbcow):

hmm 1- means approach from left ... 0.99 or something like that right by plugging in this value for x ...(1-.99) = .01 which is positive

OpenStudy (anonymous):

oh i get it

OpenStudy (dumbcow):

awesome :) sorry i couldn't explain it better...absolute value can really throw people off

OpenStudy (anonymous):

so everythign in absolute must be positive regardless of 1-. you have to change it so it becomes positive. i was thinking 1- would be negative overall and 1+ would be positive

OpenStudy (dumbcow):

no not at all , the plus and minus have nothing to do with sign , just whether you are approaching from right or left

OpenStudy (dumbcow):

obviously when x=1 then 1-x = 0 but what if x is 0.99 or 1.01 then you look at whether (1-x) is pos or neg

OpenStudy (anonymous):

for 1- i was using numbers such as -2. that would change the outcome. so how did you determine it was from 0<x<1

OpenStudy (dumbcow):

well you want to pick a number that is really close to 1 since limit means approaching a number thats why i chose 0.99 for 1- and 1.01 for 1+

OpenStudy (anonymous):

gotcha! i will definately keep that in mind. thanks again!